OpenStudy (anonymous):

Walk me through this problem plz! Solve for c 1/k + 1/c =1

OpenStudy (phi):

one way is to "clear the denominators" multiply the equation by kc (both sides and all terms) can you do that ?

OpenStudy (anonymous):

Okay. So you get k+c=kc

OpenStudy (phi):

now to solve for c, "move" the c from the left side to the right side we do this by add -c to both sides k+c-c = kc - c and simplify

OpenStudy (anonymous):

Okay so k=kc-c

OpenStudy (phi):

now the "hard part" (though not that hard) factor c out of each term in kc - c can you do that ?

OpenStudy (anonymous):

Oh i see now. k=c(k-1)

OpenStudy (phi):

finally, divide both sides by k-1

OpenStudy (anonymous):

k/k-1=c

OpenStudy (anonymous):

Does the k cancel out?

OpenStudy (phi):

yes, but I would put parens around k-1 you are dividing by k-1 \[ c = \frac{k}{k-1} \] if you write it c = k/(k-1) we need the parens otherwise it looks like c = k/k - 1 which means 1 -1 = 0 (that is a bozo no-no)

OpenStudy (anonymous):

Oh I see. Thank you.

OpenStudy (phi):

yw